OpenStudy (zeesbrat3):

Find the x-coordinates of any relative extrema and inflection point(s) for the function f of x equals 9 times x raised to the one third power plus 9 halves times x raised to the four thirds power . You must justify your answer using an analysis of f ′(x) and f ′′(x).

OpenStudy (zeesbrat3):

\[f(x) = 9x^\frac{ 1 }{ 3 } + \frac{ 9 }{ 2 }x^\frac{ 4 }{ 3 }\]

OpenStudy (zeesbrat3):

I have that \[f'(x) = 3x^\frac{ 1 }{ 3 }[x ^{-2} + 2]\]

OpenStudy (zeesbrat3):

@ganeshie8

OpenStudy (zeesbrat3):

So, I set f'(x) = 0 and I got \[x = 0, \pm \frac{ 1 }{ \sqrt{-2} }\]

OpenStudy (zeesbrat3):

I'm confused...

OpenStudy (zeesbrat3):

@jim_thompson5910 @amistre64

OpenStudy (zeesbrat3):

Hi, Jim_thomspon5910

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = 9x^{1/3} + \frac{9}{2}x^{4/3}\] \[\Large f \ '(x) = 9(1/3)x^{1/3-1} + \frac{9}{2}*\frac{4}{3}x^{4/3-1}\] \[\Large f \ '(x) = 3x^{-2/3} + 6x^{1/3}\] \[\Large f \ '(x) = 3x^{1/3}\left(x^{-2} + 2\right)\] so I'm getting the same f ' (x) as you did

OpenStudy (zeesbrat3):

Then set it to 0, yes?

jimthompson5910 (jim_thompson5910):

if you set f ' (x) equal to zero, and solve for x, you do get \[\Large x = 0 \text{ or } x = \pm \frac{1}{\sqrt{-2}}\] so I agree there as well

jimthompson5910 (jim_thompson5910):

The portion \(\Large \pm\frac{1}{\sqrt{-2}}\) aren't real numbers, so we can ignore these two solutions

jimthompson5910 (jim_thompson5910):

Now either do a first derivative test or a second derivative test to see if a local min or max is at x = 0

OpenStudy (zeesbrat3):

The number line check?

jimthompson5910 (jim_thompson5910):

yeah you can do that for a first derivative test

jimthompson5910 (jim_thompson5910):

see if f ' (x) changes sign for a value less than 0, and for a value greater than 0

OpenStudy (zeesbrat3):

when it is less than 0, it is decreasing, when it is greater than 0, it is increasing

jimthompson5910 (jim_thompson5910):

correct on both

OpenStudy (zeesbrat3):

So 0 is a min

jimthompson5910 (jim_thompson5910):

when x < 0, f ' (x) < 0 when x > 0, f ' (x) > 0 we have a local min at x = 0 

jimthompson5910 (jim_thompson5910):

the min is at x = 0 the actual min value itself is unknown at this point

jimthompson5910 (jim_thompson5910):

plug x = 0 back into f(x) to find the min value

jimthompson5910 (jim_thompson5910):

I'm not thinking if x = 0, then f ' (x) is undefined because of the x^(-2) portion. That leads to a division by zero error. So x = 0 cannot be a critical value. The graph of f(x) also shows a vertical tangent at x = 0

OpenStudy (zeesbrat3):

Why plug 0 back into f(x)?

ganeshie8 (ganeshie8):

\[\Large f(x) = 9x^{1/3} + \frac{9}{2}x^{4/3}\] \[\Large f \ '(x) = 9(1/3)x^{1/3-1} + \frac{9}{2}*\frac{4}{3}x^{4/3-1}\] \[\Large f \ '(x) = 3x^{-2/3} + 6x^{1/3}\] \[\Large f \ '(x) = 3x^{1/3}\left(x^{-\color{red}{1}} + 2\right)\]

jimthompson5910 (jim_thompson5910):

oh my bad, I factored incorrectly